# Data-informativity conditions for structured linear systems with implications for dynamic networks

### v3

@@ Abstract
thanks: Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the European Research Council can be held responsible for them.thanks: This work was supported by VINNOVA Competence Center AdBIOPRO, contract [2016-05181] and by the Swedish Research Council through the research environment NewLEADS (New Directions in Learning Dynamical Systems), contract [2016-06079], and contract [2019-04956].thanks: Original version 20 November 2024; revised 29 July 2025.
@@ 1 Introduction
While the dynamic systems that are currently addressed in engineering and science, become more and more complex, large-scale and interconnected, the attention for modelling, monitoring and control aspects of interconnected systems, like e.g. dynamic networks, is growing. For the resulting control problems this has led to the development of decentralized, distributed and multi-agent type of control systems. On the modelling side, the development of analysis tools for systems over graphs has to be mentioned, and we have seen the development of a field called data-driven modeling (or system identification) in dynamic networks. This paper is situated in this latter area.
Identification of/in dynamic networks has been developed in a framework where dynamic systems are interconnected in a particular interconnection structure (topology), including cycles, that may be known a priori. In such a setting several different problems have been considered. Following the early contributions on uniqueness of network representations Goncalves&Warnick:08 and topology estimation Materassi&Innocenti:10 ; Materassi&Salapaka:12 , a framework for network identification in a prediction error setting is introduced in VandenHof&etal_Autom:13 , while subspace methods for network identification have been developed in Haber&Verhaegen_TAC:14 ; Verhaegen&etal:22 .
@@
The question that we address in this paper is therefore: can we exploit structural information (topology) of the network, in the form of a priori known entries in the predictor model, to relax the data-informativity conditions and to focus these conditions on estimating a single target module only. In comparison with the current literature, these are two innovative steps to be taken: including the structural information, and addressing a single target module only.
@@ 4 Data-informativity conditions for consistent estimation
$\varepsilon(t,\theta)$ $=$ $H(q,\theta)^{-1}[G{\color[rgb]{0,0,0}{}^{0}}(q)-G(q,\theta)]u(t)+$ (7)
@@ 5.2 Including a network structure in the generation of input signals
$u(t)={\color[rgb]{0,0,0}P}(q)u(t)+{\color[rgb]{0,0,0}Q}(q)x(t)$ (13)
with $P$ a $m\times m$ proper and stable transfer function matrices, with (IP(q))(I-P(q)) invertible, $P$ being hollow, meaning that ${\color[rgb]{0,0,0}P(q)}]_{kk}=0$ for all $k$, and $\color[rgb]{0,0,0}Q$} being composed of a (sub)set of columns of the $m\times m$ identity matrix. Furthermore, and without loss of generality, it is assumed that $\Phi_{x}(\omega)$ is diagonal and $\Phi_{x}(\omega)\succ 0$ for all $\omega$. An example of a network graph that generates an input vector $u$ is shown in Figure 2, in the particular situation of a target module $G_{j1}$.
@@
  • $(u_{k},u_{\ell})\in\mathcal{E}$ if and only if ${\color[rgb]{0,0,0}P}]_{\ell k}\neq 0$;
  • $(x_{k},u_{\ell})\in\mathcal{E}$ if and only if ${\color[rgb]{0,0,0}Q}]_{\ell k}\neq 0$.
@@
A vertex set $\color[rgb]{0,0,0}\mathcal{V}_{\mathcal{D}^{c}}$} is called a disconnecting set from $\mathcal{V}_{1}$ to $\mathcal{V}_{2}$, if upon removal of $\color[rgb]{0,0,0}\mathcal{V}_{\mathcal{D}^{c}}$} from the graph, there is no directed path from $\mathcal{V}_{1}$ to $\mathcal{V}_{2}$. It is a minimum disconnecting set if it has the minimum cardinality among all such disconnecting sets. $\Box$
@@
As a next step we are going to decompose the input signals in $\kappa_{u}^{[j]}$ in three different components. In doing so we adopt the notation that for an $m$-dimensional node vector $u$, and $\mathcal{D}^{c}\subset[1,\cdots m]$ an index set, $u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ is a vector of those components of $u$ that are defined by the index set $\mathcal{D}^{c}$. We will use this general notation throughout the remaining part of the paper.
@@
The interpretation of this disconnecting set is that it disconnects the input $u_{i}$ of the target module from all other inputs to parametrized modules. To illustrate this, consider the situation as depicted in Figure 2. The disconnecting set $\{u_{2}\}$ disconnects ${\color[rgb]{0,0,0}\{}u_{1}{\color[rgb]{0,0,0}\}}$ from ${\color[rgb]{0,0,0}\{}u_{3}{\color[rgb]{0,0,0}\}}$. As a result, the dependency of $u_{3}$ on $u_{1}$ is fully covered by the signal $u_{2}$. This phenomenon is closely related to the concept of blocked parallel paths from $u_{1}$ to $y_{j}$, i.e. parallel to the target module $G_{j1}$ (Dankers&etal_TAC:16, ), i.e. $u_{2}$ blocks all connections from $u_{1}$ to $y_{j}$ that are not passing through $G_{j1}$.
@@
  • Let $u_{(i\cup\!{\scriptscriptstyle\mathcal{D}}^{c})\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}$ be the projection [^2] Projection is considered with respect to the inner product defined by $x,y\!>:=\bar{\mathbb{E}}[x(t)y^{T}(t)]$. of signals $u_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}$ onto the orthogonal complement of $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$.
@@ 6 Example
First we consider the situation that all four modules $G_{jk}(q,\theta)$, $k=1,\cdots 4$ are parametrized. A minimum disconnecting set ${\color[rgb]{0,0,0}\{}u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}{\color[rgb]{0,0,0}\}}$ from input $u_{i}=u_{1}$ to all other inputs of parametrized modules is ${\color[rgb]{0,0,0}\{}u_{2}{\color[rgb]{0,0,0}\}}$. In an alternative interpretation it is clear that $u_{2}$ blocks all parallel paths (i.e. parallel to the target module $G_{j1}$) from $u_{i}$ to output $y_{j}$.
The signal $u_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}$ is composed of $u_{i}=u_{1}$, $u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}=u_{2}$, while $u_{\!{\scriptscriptstyle\mathcal{T}}}=(u_{3},u_{4}){\color[rgb]{0,0,0}{}^{T}}$. Possible excitation or noise signals $x_{3}$ and $x_{4}$ will serve as $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$. Since $u$ and $e$ are considered to be independent, and $e$ is white noise, the condition (16) can be replaced by a positive definite spectrum of $u_{(1\cup 2)\perp(x_{3},x_{4})}$.
@@
$T_{1}(q,\theta)$ $=$ $\Delta G_{j1}+\Delta G_{j2}\frac{{\color[rgb]{0,0,0}P}_{21}}{1\!-\!{\color[rgb]{0,0,0}P}_{32}{\color[rgb]{0,0,0}P}_{23}}\!+\!\Delta G_{j3}\frac{{\color[rgb]{0,0,0}P}_{32}{\color[rgb]{0,0,0}P}_{21}}{1\!-\!{\color[rgb]{0,0,0}P}_{32}{\color[rgb]{0,0,0}P}_{23}}$
$T_{2}(q,\theta)$ $=$ $\Delta G_{j2}\frac{1}{1-{\color[rgb]{0,0,0}P}_{32}{\color[rgb]{0,0,0}P}_{23}}+\Delta G_{j3}\frac{{\color[rgb]{0,0,0}P}_{32}}{1-{\color[rgb]{0,0,0}P}_{32}{\color[rgb]{0,0,0}P}_{23}}$
$T_{3}(q,\theta)$ $=$ $\Delta G_{j2}\frac{{\color[rgb]{0,0,0}P}_{23}}{1-{\color[rgb]{0,0,0}P}_{32}{\color[rgb]{0,0,0}P}_{23}}+\Delta G_{j3}\frac{1}{1-{\color[rgb]{0,0,0}P}_{32}{\color[rgb]{0,0,0}P}_{23}}$
$T_{4}(q,\theta)$ $=$ $\Delta G_{j2}\frac{{\color[rgb]{0,0,0}P}_{23}{\color[rgb]{0,0,0}P}_{34}}{1\!-\!{\color[rgb]{0,0,0}P}_{32}{\color[rgb]{0,0,0}P}_{23}}\!+\!\Delta G_{j3}\frac{{\color[rgb]{0,0,0}P}_{34}}{1\!-\!{\color[rgb]{0,0,0}P}_{32}{\color[rgb]{0,0,0}P}_{23}}\!+\!\Delta G_{j4}$
where $\Delta G_{jk}:=G_{jk}^{0}(q)-G_{jk}(q,\theta)$, for $k=1,\cdots 4$. The minimum of the cost function $\bar{\mathbb{E}}\varepsilon^{T}(t,\theta)\varepsilon(t,\theta)$ is achieved if the $T_{k}$-dependent terms are zero. Presence of $x_{k}$ implies that $T_{k}(q,\theta)=0$ in the minimum of the cost function. If $x_{1}$ and $x_{2}$ are present and persistently exciting, then $T_{1}$ and $T_{2}$ are forced to be zero, which implies that $\Delta G_{j1}=0$, and thus a consistent estimate of $G^{0}_{j_{1}}$ is guaranteed. This follows from the observation that $T_{1}(q,\theta)=\Delta G_{j1}+{\color[rgb]{0,0,0}P}_{21}T_{2}(q,\theta)$. If we would replace $x_{2}$ by presence of $x_{3}$, a similar result does not exist. Introducing $x_{4}$ shows that $T_{4}(q,\theta)$ gets involved in the expression and a new unknown term $\Delta G_{j4}$ is introduced also. This explains that forcing $T_{1}$ and $T_{4}$ to be zero is not sufficient for arriving at $\Delta G_{j1}=0$, or in other words: exciting $x_{1}$ and $x_{4}$ does not lead to a consistent estimate of the target module.
In the case that module $G_{j2}(q)$ is not parametrized, but known a priori, there are two options for choosing minimum disconnecting sets:
  • When selecting $u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}=u_{2}$, it follows that $u_{\!{\scriptscriptstyle\mathcal{T}}}=(u_{3},u_{4}){\color[rgb]{0,0,0}{}^{T}}$, and as a result $x_{3}$ and $x_{4}$ cannot be used for excitation of $(u_{1},u_{2})$, and therefore we need excitations $x_{1}$ and $x_{2}$.
@@
In the case that $G_{j4}$ is not parametrized and known a priori, the minimum disconnecting set from $u_{1}$ to ${\color[rgb]{0,0,0}\{}u_{2},u_{3}{\color[rgb]{0,0,0}\}}$ is given by $u_{2}$. Then $u_{\!{\scriptscriptstyle\mathcal{T}}}=u_{3}$, while the external signals $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$ that affect $u_{\!{\scriptscriptstyle\mathcal{T}}}$ are given by $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}=(x_{3},x_{4}){\color[rgb]{0,0,0}{}^{T}}$. This implies that (only) signals $(x_{1},x_{2})$ are available for realizing a positive definite spectrum of $(u_{1},u_{2})_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}$. Verification of this through the expressions of $T_{k}(q,\theta)$, shows that with $\Delta G_{j4}=0$, the linear combination of the two terms in $T_{3}$ and $T_{4}$ are different from the linear combinations of $\Delta G_{j2}$ and $\Delta G_{j3}$ in the expressions for $T_{1}$ and $T_{2}$. As a result the original result with all modules parametrized follows as the solution.
@@ 7 Path-based interpretation
The spectrum condition $\Phi_{\mu^{[j]}}\succ 0$ for almost all $\omega$, holds generically [^3] In line with its use in Hendrickx&Gevers&Bazanella_TAC:19 , this concerns a property that holds for all possible dynamic modules in the network, except possibly for a set of measure 0.if in the graph of the network there exist $dim(\mu^{[j]})$ vertex disjoint paths from the node sets ${\color[rgb]{0,0,0}\{}x,e{\color[rgb]{0,0,0}\}}$ to ${\color[rgb]{0,0,0}\{}\mu^{[j]}{\color[rgb]{0,0,0}\}}$, and the signals in $x$ are mutually independent and persistently exciting.
@@
The graph-based condition in Proposition 5 is satisfied if there exist $dim(u_{\!{\scriptscriptstyle\mathcal{D}}^{c}})+1$ vertex disjoint paths between the node sets $x\backslash x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}\}$ and ${\color[rgb]{0,0,0}\{}u_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}{\color[rgb]{0,0,0}\}}$.
@@
We can illustrate this result for the example in Section 6. In the situation where all $G$-modules are parametrized it was found that ${\color[rgb]{0,0,0}\{}u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}{\color[rgb]{0,0,0}\}}=\{u_{2}\}$, ${\color[rgb]{0,0,0}\{}u_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}{\color[rgb]{0,0,0}\}}=\{u_{1},u_{2}\}$ and $x\backslash x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}\}=\{x_{1},x_{2}\}$. Therefore, two vertex disjoint paths between $\{x_{1},x_{2}\}$ and $\{u_{1},u_{2}\}$, are sufficient to guarantee (generic) data-informativity. This condition is satisfied if both $x_{1}$ and $x_{2}$ are present.

### v4

@@ Abstract
thanks: Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the European Research Council can be held responsible for them.thanks: This work was supported by VINNOVA Competence Center AdBIOPRO, contract [2016-05181] and by the Swedish Research Council through the research environment NewLEADS (New Directions in Learning Dynamical Systems), contract [2016-06079], and contract [2019-04956].thanks: Original version 20 November 2024; revised 29 July 2025, 7 January 2026.
@@ 1 Introduction
While the dynamic systems that are currently addressed in engineering and science, become more and more complex, large-scale and interconnected, the attention for modelling, monitoring and control aspects of interconnected systems, like e.g. dynamic networks, is growing. For the resulting control problems this has led to the development of decentralized, distributed and multi-agent type of control systems. On the modelling side, the development of analysis tools for systems over graphs has to be mentioned, and we have seen the development of a field called data-driven modeling (or system identification) in dynamic networks. This paper is situated in this latter area, that will be overviewed next.
Identification of/in dynamic networks has been developed in a framework where dynamic systems are interconnected in a particular interconnection structure (topology), including cycles, that may be known a priori. In such a setting several different problems have been considered. Following the early contributions on uniqueness of network representations Goncalves&Warnick:08 and topology estimation Materassi&Innocenti:10 ; Materassi&Salapaka:12 , a framework for network identification in a prediction error setting is introduced in VandenHof&etal_Autom:13 , while subspace methods for network identification have been developed in Haber&Verhaegen_TAC:14 ; Verhaegen&etal:22 . Alternatively, there are network identification approaches based on conditional independence and dynamic factor models, see e.g. Zorzi&Chiuso:17 ; Veedu&Doddi&Salapaka_TAC:22 .
@@
The question that we address in this paper is therefore: in the setting of the mentioned local direct method Ramaswamy&VandenHof_TAC:21 , can we exploit structural information (topology) of the network, in the form of a priori known entries in the predictor model, to relax the data-informativity conditions and to focus these conditions on estimating a single target module only. In comparison with the current data-informativity literature, these are two innovative steps to be taken: including the structural information, and addressing a single target module only.
@@ 4 Data-informativity conditions for consistent estimation
$\varepsilon(t,\theta)$ $=$ $H(q,\theta)^{-1}[G^{0}(q)-G(q,\theta)]u(t)+$ (7)
@@ 5.2 Including a network structure in the generation of input signals
$u(t)=P(q)u(t)+Q(q)x(t)$ (13)
with $P$ a $m\times m$ proper and stable transfer function matrices, with (IP(q))(I-P(q)) invertible, $P$ being hollow, meaning that $P(q)]_{kk}=0$ for all $k$, and $Q$ being composed of a (sub)set of columns of the $m\times m$ identity matrix. Furthermore, and without loss of generality, it is assumed that $\Phi_{x}(\omega)$ is diagonal and $\Phi_{x}(\omega)\succ 0$ for all $\omega$. An example of a network graph that generates an input vector $u$ is shown in Figure 2, in the particular situation of a target module $G_{j1}$.
@@
  • $(u_{k},u_{\ell})\in\mathcal{E}$ if and only if $P]_{\ell k}\neq 0$;
  • $(x_{k},u_{\ell})\in\mathcal{E}$ if and only if $Q]_{\ell k}\neq 0$.
@@
A vertex set $\mathcal{V}_{\mathcal{D}^{c}}$ is called a disconnecting set from $\mathcal{V}_{1}$ to $\mathcal{V}_{2}$, if upon removal of $\mathcal{V}_{\mathcal{D}^{c}}$ from the graph, there is no directed path from $\mathcal{V}_{1}$ to $\mathcal{V}_{2}$. It is a minimum disconnecting set if it has the minimum cardinality among all such disconnecting sets. $\Box$
Note that $\mathcal{V}_{\mathcal{D}^{c}}$ may contain elements of $\mathcal{V}_{1}$ and/or $\mathcal{V}_{2}$.
@@
As a next step we are going to decompose the input signals in $\kappa_{u}^{[j]}$ in three different components. In doing so we adopt the notation that for an $m$-dimensional node vector $u$, and $\mathcal{D}^{c}\subset{\color[rgb]{0,0,0}\{}1,\cdots m{\color[rgb]{0,0,0}\}}$ an index set, $u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ is a vector of those components of $u$ that are defined by the index set $\mathcal{D}^{c}$. We will use this general notation throughout the remaining part of the paper.
@@
The interpretation of this disconnecting set is that it disconnects the input $u_{i}$ of the target module from all other inputs to parametrized modules. To illustrate this, consider the situation as depicted in Figure 2. The disconnecting set $\{u_{2}\}$ disconnects $\{u_{1}\}$ from $\{u_{3}\}$. As a result, the dependency of $u_{3}$ on $u_{1}$ is fully covered by the signal $u_{2}$. This phenomenon is closely related to the concept of blocked parallel paths from $u_{1}$ to $y_{j}$, i.e. parallel to the target module $G_{j1}$ (Dankers&etal_TAC:16, ), i.e. $u_{2}$ blocks all connections from $u_{1}$ to $y_{j}$ that are not passing through $G_{j1}$.
@@
  • Let $u_{(i\cup\!{\scriptscriptstyle\mathcal{D}}^{c})\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}$ be the projection [^2] Projection is considered with respect to the inner product defined by $x_{m},y_{n}\!>:=\bar{\mathbb{E}}[x_{m}(t)y_{n}(t)]$, while the projection of a vector signal $x$ onto a vector signal $y$ is defined as the projection of every component of vector $x$ onto all components of vector $y$. of signals $u_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}$ onto the orthogonal complement of $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$.
@@ 6 Example
First we consider the situation that all four modules $G_{jk}(q,\theta)$, $k=1,\cdots 4$ are parametrized. A minimum disconnecting set $\{u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\}$ from input $u_{i}=u_{1}$ to all other inputs of parametrized modules is $\{u_{2}\}$. In an alternative interpretation it is clear that $u_{2}$ blocks all parallel paths (i.e. parallel to the target module $G_{j1}$) from $u_{i}$ to output $y_{j}$.
The signal $u_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}$ is composed of $u_{i}=u_{1}$, $u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}=u_{2}$, while $u_{\!{\scriptscriptstyle\mathcal{T}}}=(u_{3},u_{4})^{T}$. Possible excitation or noise signals $x_{3}$ and $x_{4}$ will serve as $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$. Since $u$ and $e$ are considered to be independent, and $e$ is white noise, the condition (16) can be replaced by a positive definite spectrum of $u_{(1\cup 2)\perp(x_{3},x_{4})}$.
@@
$T_{1}(q,\theta)$ $=$ $\Delta G_{j1}+\Delta G_{j2}\frac{P_{21}}{1\!-\!P_{32}P_{23}}\!+\!\Delta G_{j3}\frac{P_{32}P_{21}}{1\!-\!P_{32}P_{23}}$
$T_{2}(q,\theta)$ $=$ $\Delta G_{j2}\frac{1}{1-P_{32}P_{23}}+\Delta G_{j3}\frac{P_{32}}{1-P_{32}P_{23}}$
$T_{3}(q,\theta)$ $=$ $\Delta G_{j2}\frac{P_{23}}{1-P_{32}P_{23}}+\Delta G_{j3}\frac{1}{1-P_{32}P_{23}}$
$T_{4}(q,\theta)$ $=$ $\Delta G_{j2}\frac{P_{23}P_{34}}{1\!-\!P_{32}P_{23}}\!+\!\Delta G_{j3}\frac{P_{34}}{1\!-\!P_{32}P_{23}}\!+\!\Delta G_{j4}$
where $\Delta G_{jk}:=G_{jk}^{0}(q)-G_{jk}(q,\theta)$, for $k=1,\cdots 4$. The minimum of the cost function $\bar{\mathbb{E}}\varepsilon^{T}(t,\theta)\varepsilon(t,\theta)$ is achieved if the $T_{k}$-dependent terms are zero. Presence of $x_{k}$ implies that $T_{k}(q,\theta)=0$ in the minimum of the cost function. If $x_{1}$ and $x_{2}$ are present and persistently exciting, then $T_{1}$ and $T_{2}$ are forced to be zero, which implies that $\Delta G_{j1}=0$, and thus a consistent estimate of $G^{0}_{j_{1}}$ is guaranteed. This follows from the observation that $T_{1}(q,\theta)=\Delta G_{j1}+P_{21}T_{2}(q,\theta)$. If we would replace $x_{2}$ by presence of $x_{3}$, a similar result does not exist. Introducing $x_{4}$ shows that $T_{4}(q,\theta)$ gets involved in the expression and a new unknown term $\Delta G_{j4}$ is introduced also. This explains that forcing $T_{1}$ and $T_{4}$ to be zero is not sufficient for arriving at $\Delta G_{j1}=0$, or in other words: exciting $x_{1}$ and $x_{4}$ does not lead to a consistent estimate of the target module.
In the case that module $G_{j2}(q)$ is not parametrized, but known a priori, there are two options for choosing (minimum) disconnecting sets:
  • When selecting $u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}=u_{2}$, it follows that $u_{\!{\scriptscriptstyle\mathcal{T}}}=(u_{3},u_{4})^{T}$, and as a result $x_{3}$ and $x_{4}$ cannot be used for excitation of $(u_{1},u_{2})$, and therefore we need excitations $x_{1}$ and $x_{2}$.
@@
In the case that $G_{j4}$ is not parametrized and known a priori, the (minimum{)} disconnecting set from $u_{1}$ to $\{u_{2},u_{3}\}$ is given by $u_{2}$. Then $u_{\!{\scriptscriptstyle\mathcal{T}}}=u_{3}$, while the external signals $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$ that affect $u_{\!{\scriptscriptstyle\mathcal{T}}}$ are given by $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}=(x_{3},x_{4})^{T}$. This implies that (only) signals $(x_{1},x_{2})$ are available for realizing a positive definite spectrum of $(u_{1},u_{2})_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}$. Verification of this through the expressions of $T_{k}(q,\theta)$, shows that with $\Delta G_{j4}=0$, the linear combination of the two terms in $T_{3}$ and $T_{4}$ are different from the linear combinations of $\Delta G_{j2}$ and $\Delta G_{j3}$ in the expressions for $T_{1}$ and $T_{2}$. As a result the original result with all modules parametrized follows as the solution.
@@ 7 Path-based interpretation
The spectrum condition $\Phi_{\mu^{[j]}}\succ 0$ for almost all $\omega$, holds generically [^3] In line with its use in Hendrickx&Gevers&Bazanella_TAC:19 , this concerns a property that holds for all possible dynamic modules in the network, except possibly for a set of measure 0. if in the graph of the network there exist $dim(\mu^{[j]})$ vertex disjoint paths from the node sets $\{x,e\}$ to $\{\mu^{[j]}\}$, and the signals in $x$ are mutually independent and persistently exciting.
@@
The graph-based condition in Proposition 5 is satisfied if there exist $dim(u_{\!{\scriptscriptstyle\mathcal{D}}^{c}})+1$ vertex disjoint paths between the node sets $x\backslash x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}\}$ and $\{u_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}\}$.
@@
We can illustrate this result for the example in Section 6. In the situation where all $G$-modules are parametrized it was found that $\{u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\}=\{u_{2}\}$, $\{u_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}\}=\{u_{1},u_{2}\}$ and $x\backslash x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}\}=\{x_{1},x_{2}\}$. Therefore, two vertex disjoint paths between $\{x_{1},x_{2}\}$ and $\{u_{1},u_{2}\}$, are sufficient to guarantee (generic) data-informativity. This condition is satisfied if both $x_{1}$ and $x_{2}$ are present.